### v3

# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

Abstract
This article proves the following theorem, first enunciated by Roger Penrose about 70 years ago: In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). This conic is unique if not all the common points of condition (ii) are the same. The proof is based on the following analogous theorem, which is also proved: In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists an eighth quadric such that the completed cube satisfies (i) and (ii). This quadric is unique if not all the common axes of condition (ii) are the same.
1 Introduction
Let us start with a classical theorem from 340 AD by Pappus of Alexandria (290–350 AD) [Jon13] and its dual counterpart (Figure 1).
Consider a pair of lines in the plane (call them $\ell$ and $\ell^{\prime}$). Assign three arbitrary points on each line ($A,B,C$ on $\ell$ and $A^{\prime},B^{\prime},C^{\prime}$ on $\ell^{\prime}$). Join these three pairs of points in a crossing manner, giving us three meeting points (namely the meet of lines $AB^{\prime}$ and $A^{\prime}B$, the meet of lines $BC^{\prime}$ and $B^{\prime}C$, and the meet of lines $CA^{\prime}$ and $C^{\prime}A$). Then these three points must be collinear.
Consider a pair of points in the plane (call them $L$ and $L^{\prime}$). Draw three lines through each point ($a,b,c$ through $L$ and $a^{\prime},b^{\prime},c^{\prime}$ through $L^{\prime}$). Meet these three pairs of lines in a diagonal manner, which are joined into three lines (namely the join of points $ab^{\prime}$ and $a^{\prime}b$, the join of points $bc^{\prime}$ and $b^{\prime}c$, and the join of points $ca^{\prime}$ and $c^{\prime}a$). Then these three lines must concur.
[…]
Figure 1: Pappus’ theorem (340 AD) (left) and its dual theorem (right).
A fascinating quality about Pappus’ theorem is that the statement of the theorem requires very few structures about the plane, that two distinct points can be joined into a unique line, and two distinct lines meet at a unique point (possibly at the line at infinity), like those rules in perspective drawings. No lengths, angles, or parallelisms are involved. This is a projective geometric theorem. Note that one may swap “points” and ”lines” and swap “meet” and “join” in a planar projective geometric theorem to obtain the dual theorem, as shown above.
It is fair to say that neither Pappus’ theorem and its dual are obvious. A natural question is whether there is a general pattern to theorems like Pappus’, so that upon this understanding, Pappus’ theorem becomes more intuitive and easier to discover.
Blaise Pascal (1623–1662), at age 16, discovered a remarkable generalization of Pappus’ theorem, known as Pascal’s mystic hexagram [DC20]. The initial line pair of Pappus’ theorem, as well as the initial point pair of the dual Pappus theorem, is a degeneration of a conic section. After the line pair is replaced by an arbitrary conic, the rest of the theorem statement remains true (Figure 2).
[…]
Figure 2: Pascal’s theorem (1640) (left) and its dual theorem—Brianchon’s theorem (1807) (right).
Consider a conic in the plane. Assign six points on the conic (labeled $A,B,C,A^{\prime},B^{\prime},C^{\prime}$). Join these three pairs of points in a crossing manner, giving us three meeting points (namely the meet of lines $AB^{\prime}$ and $A^{\prime}B$, the meet of lines $BC^{\prime}$ and $B^{\prime}C$, and the meet of lines $CA^{\prime}$ and $C^{\prime}A$). Then these three points must be collinear.
Consider a conic in the plane. Draw six tangent lines of the conic (labeled $a,b,c,a^{\prime},b^{\prime},c^{\prime}$). Meet these three pairs of lines in a diagonal manner, which are joined into three lines (namely the join of points $ab^{\prime}$ and $a^{\prime}b$, the join of points $bc^{\prime}$ and $b^{\prime}c$, and the join of points $ca^{\prime}$ and $c^{\prime}a$). Then these three lines must concur.
The dual theorem (right-hand side) is known as Brianchon’s Theorem, named after Charles Julien Brianchon (1783–1864).
In 1848, George Salmon published a textbook on conic sections, which include a further generalization of the Pascal–Brianchon theorem [Sal17, §264], rediscovered by Evelyn, Money-Coutts, and Tyrrell in 1974 [EMCT74, §2.3]. The three point pairs $AA^{\prime}$, $BB^{\prime}$, $CC^{\prime}$ of Pascal theorem are degeneration of thin ellipses that are in double contact with the initial conic. And the lines $AB^{\prime}$ and $A^{\prime}B$ are the limit of a pair of common tangents between the two thin ellipses collapsing to $AA^{\prime}$ and $BB^{\prime}$. After replacing the three point pairs by three conics in double contact with a fourth, the remainder of Pappus’ theorem statement still holds (Figure 3).
[…]
Figure 3: Salmon’s theorem (1848) and its dual theorem.
If three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ have each in double contact with a fourth $\mathbf{{S}}_{0}$, then the three intersection points, each from a pair of common tangents between $\mathbf{{T}}^{1}\mathbf{{T}}^{2}$, between $\mathbf{{T}}^{2}\mathbf{{T}}^{3}$, and between $\mathbf{{T}}^{3}\mathbf{{T}}^{1}$, are collinear.
If three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ have each in double contact with a fourth $\mathbf{{S}}_{0}$, then a common chord between $\mathbf{{T}}^{1}\mathbf{{T}}^{2}$, a common chord between $\mathbf{{T}}^{2}\mathbf{{T}}^{3}$, and a common chord between $\mathbf{{T}}^{3}\mathbf{{T}}^{1}$, are concurrent. [^6] The original statement is “If three conics have each in double contact with a fourth, then six of their chords of intersection form a complete quadrilateral.”
The reader may have observed a pattern of this progression of generalizing Pappus’ theorem. A line pair and a point pair are replaced by conic sections. The incidence relations of point pairs on a conic or line pairs tangent to a conic are both unified by two conics in double contact.
Refer to captionRefer to caption
Figure 4: The main theorem—the eight conic theorem—first enunciated by Roger Penrose in around 1950. The right figure illustrates an instance of the configuration where all 8 conics are ellipses, shown next to the corresponding cube graph of conics and double contact relationships.
This article shows a “final” generalization of this progression of projective geometric theorems, first stated by Roger Penrose, the last author of this article, in around 1950 during his college years. We refer to it Penrose’s eight-conic theorem (Figure 4):
If conics are assigned to seven of the vertices of a combinatorial cube such that conics connected by an edge are in double contact, then (upon a mild consistency condition on the cube face) there exists an eighth conic which completes the cube.
Namely, if three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ each in double contact with a fourth $\mathbf{{S}}_{0}$, and another three conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ are respectively in double contact with $\mathbf{{T}}^{2},\mathbf{{T}}^{3}$, with $\mathbf{{T}}^{3},\mathbf{{T}}^{1}$, and with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$ (chosen in a consistent branch of a double contact family), then there exists a conic $\mathbf{{T}}^{0}$ that has double contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$.
Here, the consistency condition of the cube face is that the four chords of contact from the four conics have a common point. This condition is automatically guaranteed if one picks a consistent branch of double contact conics with two given conic (out of 3 branches).
This article provides a completion of one of Roger Penrose’s proof ideas for the eight-conic theorem. This idea emerged during his graduate studies at Cambridge, which is that the eight-conic configuration is a planar slice of eight quadrics in ring contact in 3D. (Two quadrics have ring contact if they touch along a conic.) It turns out that the following Penrose’s eight-quadric theorem is much more intuitive to see how it has to hold, as illustrated later in this article:
If quadrics in 3D are assigned to seven of the vertices of a combinatorial cube such that quadrics connected by an edge are in ring contact, then (upon a mild consistency condition on the cube face) there exists an eighth quadric which completes the cube.
The face consistency condition is that the four planes containing the four rings of the ring contacts must meet along a common axis. Unlike in 2D, there are generally no incompatible quadrics in 3D when completing the vertices of the cube, except in very special situations.
The article is structured as follows. Section 2 provides the necessary background on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.3, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 4. Section 5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical theorems in the projective plane. The final sections of the article provide references to related works and future outlooks.
2 Background and notation
2.1 Conics and the projective plane
We recall that a point $P$ in the projective plane $\mathbb{R}P^{2}$ can be represented by an equivalence class of vectors in a 3 dimensional real [^7] We work primarily over the real field but allow complex coefficients when they arise through algebraic closure. vector space $\mathtt{V}$:
$P=\{cV:c\in\mathbb{R}\setminus\{0\}\}$ (1)
for $V\in\mathtt{V}\setminus\{0\}$. A line $p$ can be described as its set of incident points by means of a vector $v\in\mathtt{V}^{*}\setminus\{0\}$ in the dual vector space of $\mathtt{V}$, the space of linear functions $\mathtt{V}\rightarrow\mathbb{R}$:
$p=\{P\in\mathbb{R}P^{2}:v(V)=0,V\in P\}.$ (2)
The lines also naturally form a projective $2$-space called the projective dual of ${\mathbb{R}P^{2}}^{*}$ by representing them as equivalence classes of dual vectors:
$p=\{cv:c\in\mathbb{R}\setminus\{0\}\}.$ (3)
for $v\in\mathtt{V}^{*}\setminus\{0\}$. We could equally well have started by describing lines according to equation (3) and proceeded to points by describing them as the sets of incident lines:
$P=\{p\in\mathbb{R}P^{2}:v(V)=0,v\in p\}$ (4)
according to the fact that for a finite dimensional vector space, the dual space of the dual space is isomorphic to the original space according to the mapping $v\mapsto f_{V}$:
$f_{V}(v)=v(V).$
thus the roles of the two are completely interchangeable: the principle of duality. The symmetrical relationship of the two can be expressed with the notation
$\langle v,V\rangle=v(V).$
A point-line correlation is an invertible map $\mathbb{R}P^{2}\mapsto{\mathbb{R}P^{2}}^{*}$ induced by a linear transformation $\mathtt{V}\mapsto\mathtt{V}^{*}$. The fundamental theorem of projective geometry states that this is the most general type of mapping between the two which always sends any three collinear points to three copunctual lines. A line-point correlation is a similarly defined map ${\mathbb{R}P^{2}}^{*}\mapsto{\mathbb{R}P^{2}}$.
Given an ordered basis $\mathcal{B}=(V_{0},V_{1},V_{2})$ of $\mathtt{V}$, representing points $(P_{0},P_{1},P_{2})$, the dual basis $\mathcal{B}^{*}=(v_{0},v_{1},v_{2})$,
$\langle v_{j},V_{k}\rangle=\begin{cases}0&\text{if }k\neq j\\ 1&\text{if }k=j,\end{cases}$
represents the lines $(p_{0},p_{1},p_{2})$ constituting the sides of the triangle $(P_{0},P_{1},P_{2})$ with $p_{j}$ opposite to $P_{j}$. For simplicity of notation and to avoid significant repetition, we shall often identify the vectors $V$ or $v$ with the geometric elements $P$ or $p$, keeping in mind that specific representing vectors are always implied since otherwise the basis is not well defined. The point represented by $aP_{0}+bP_{1}+cP_{2}$ is often denoted by the notation $[a:b:c]$ referred to as homogeneous coordinates.
The linear transformation defining a point-line correlation can then be represented by a matrix $A$ relative to the bases $\mathcal{B}$ and $\mathcal{B}^{*}$, and vice versa for a line-point correlation. The matrix $cA$ for $c\neq 0$ clearly represents the same correlation (note however that changing the relative scaling of the vectors representing $P_{j}$ leads to a different matrix). A point-line correlation has two naturally associated line-point correlations: its inverse which has matrix $A^{-1}$ and the mapping of a line $p$ to the intersection point of any two lines mapped to by two points lying on $p$ which is well defined and has matrix $\mathrm{cof}A$, the cofactor matrix of $A$. By Cramer’s rule,
$A^{-1}=\frac{1}{\det A}\mathrm{cof}A^{T}$ (5)
where the superscript T denotes the transpose, therefore these two associated line-point correlations coincide exactly when the point-line correlation is represented by a symmetric matrix $A=A^{T}$ which is known as a polarity.
If $A$ is the matrix of a polarity, and $P$ is a point, which we identify with the column vector of a representative in the basis $(P_{0},P_{1},P_{2})$, then the line $p$ polar to $P$ is represented by a column vector, which we again identify with $p$ itself, in the basis $(p_{0},p_{1},p_{2})$:
$p=AP.$
The condition that a point is incident to its polar line may thus be expressed by the matrix equation
$p^{T}P=P^{T}AP=0.$
The locus of such points is a regular point-wise conic which is uniquely defined by the polarity and vice versa, both having a unique representative symmetric matrix (up to scaling) in a given basis. The line to which a point of a point-wise conic is both polar and incident is its tangent line. Dual to this, the locus of lines incident to their polar points is called a regular linewise conic and we refer to these points as the turning points of the linewise conic. When convenient and no ambiguity can result, we sometimes identify the matrix representing a conic in a given basis with the conic itself.
Given a regular point-wise conic, there is a uniquely associated regular linewise conic, the lines of the latter being the tangent lines of the points of the former. If the first conic is represented by the matrix $A$, the second one is represented by the inverse matrix $A^{-1}$. A degenerate point-wise conic is a conic defined in the same way but by a non-invertible symmetric matrix. In this case, three collinear points will still map to three copunctual lines when all images are defined, however, for any point $P$ whose representatives lie in the kernel of the defining matrix $A$, the image of $P$ under the degenerate polarity is geometrically undefined, since the zero vector does not represent any line in ${\mathbb{R}P^{2}}^{*}$, moreover, if $Q$ is any other point, then
$A(cP+Q)=AQ$
for any scalar $c$, so any point lying on a line through $P$ will be mapped to the same line.
  • If $\mathrm{ker}A$ has dimension $1$ then it constitutes one equivalence class in $\mathbb{R}P^{2}$ so there is a unique such point. In this case the points of a degenerate point-wise conics constitute a line pair with $P$ as their meeting point. If these two lines are called $p$ and $q$ then the matrix is given by
    $A=pq^{T}+qp^{T}$
  • If $\mathrm{ker}A$ has dimension $2$ then it constitutes the representatives of the points on a line in the projective plane. In this case the points of a degenerate point-wise conics constitute a line, referred to for algebraic reasons as a doubled line. If this line is called $p$ then the matrix is given by
    $A=pp^{T}$
Dual to this, degenerate linewise conics can be defined:
  • If $\mathrm{ker}A$ has dimension $1$ then it constitutes one equivalence class in ${\mathbb{R}P^{2}}^{*}$ so there is a unique line $p$ represented by vectors in the kernel. In this case the lines of a degenerate linewise conics constitute a point pair with $p$ as their joining line. If these to points are called $P$ and $Q$ then the matrix is given by
    $A=PQ^{T}+QP^{T}$
  • If $\mathrm{ker}A$ has dimension $2$ then it constitutes the representatives of the lines through a point in the projective plane. In this case the lines of a degenerate linewise conics constitute a point, referred to for algebraic reasons as a doubled point. If this point is called $P$ then the matrix is given by
    $A=PP^{T}$
2.1.1 Double contact
For a regular point-wise conic $A$ and a double line $pp^{T}$, the polarity defined by the conic
$B=A+cpp^{T}$ (6)
acts the same on the points of $p$, independently of the parameter $c$:
$BP=AP+cpp^{T}P=AP$ (7)
when $P$ is incident to $p$. In particular, the intersection points of $B$ with $A$ are always the intersection points of $p$ with $A$ and the tangent lines there are always the same, independently of $c$. Thus $A$ and $B$ are said to be in double contact with chord of contact $p$ and the one parameter family of conics defined by equation (6) is called a double contact pencil. The property of being in double contact is clearly self dual: two point-wise conics share tangent lines at two common points if and only if the two corresponding linewise conics share turning points at two common lines, and the fact that equation (6) is self dual, i.e., that the linewise conics of a double contact pencil can be expressed as the linear span of a regular conic and a double point: the point where the shared tangent lines meet, can be checked by taking cofactor matrices. We call this double point the apex of contact. Therefore the polarity property expressed by equation (7) can be stated as
Lemma 2.1.
The polarities on two conics in double contact act the same when restricted to a mapping between the points of the chord of contact and the lines of the apex of contact.
2.2 Quadrics and projective space
Similar definitions hold in projective 3 space $\mathbb{R}P^{3}$, using a 4 dimensional vector space $\mathtt{V}$ where now the dual space ${\mathbb{R}P^{3}}^{*}$ is the space of planes. A correlation is defined by the property that 3 collinear points map to 3 coaxial planes (three planes going through the same line: we use the word "axis" to emphasize the dual role that a line plays as the intersection of two planes in contrast to its more familiar role of connecting two points). A polarity is defined the same way but we now refer to the surface of points incident to their polar planes as a point-wise quadric. There are now 3 different cases to consider:
  • If the kernel $\mathrm{ker}A$ has dimension $1$, it represents the vertex of a cone which constitutes the points of the quadric.
  • If the kernel $\mathrm{ker}A$ has dimension $2$, it represents the axis of intersection of a plane pair. If these planes are called $p$ and $q$ then,
    $A=pq^{T}+qp^{T}.$
  • If the kernel $\mathrm{ker}A$ has dimension $3$, it is a (doubled) plane. If this plane is called $p$ then
    $A=pp^{T}$
and for plane-wise quadrics
  • If the kernel $\mathrm{ker}A$ has dimension $1$, it represents the plane of a conic curve, the planes going through a tangent line to this curve constitute the planes of the quadric.
  • If the kernel $\mathrm{ker}A$ has dimension $2$, it represents the connecting line of a point pair. If these points are called $P$ and $Q$ then,
    $A=PQ^{T}+QP^{T}.$
  • If the kernel $\mathrm{ker}A$ has dimension $3$, it is a (doubled) point. If this point is called $P$ then
    $A=PP^{T}$
The set of geometric elements represented by vectors in the kernel of $A$ is known as the degenerate set, for point-wise quadrics: the vertex of a cone, the points on the axis of a plane pair, the points on a double plane, and for plane-wise quadrics: the plane of a conic, the planes through the connecting line of a point pair or the planes through a double point. For degenerate point-wise quadrics, the polarity mapping is only defined from points, not contained in the degenerate set, to planes and vice versa for plane-wise quadrics. For point-wise quadrics:
  • The polarity on a cone maps a point which is not the vertex of the cone to a plane through the vertex, and it maps the vertex to 00 which doesn’t represent any geometric element.
  • The polarity on a plane pair maps a point not on the axis of the plane pair to a plane through the axis and any point on the axis to 00.
  • The polarity on a double plane maps any point not on the plane to the plane and any point on the plane to 00.
For plane-wise quadrics:
  • The polarity on a conic maps a plane which is not the plane of the conic to a point on the plane of the conic (in particular, the point polar to the intersection line of this plane with the plane of the conic), and it maps the plane of the conic to 00.
  • The polarity on a point pair maps a plane that does not go through the connecting line of the point pair to a point on the connecting line, and any plane through the connecting line to 00.
  • The polarity on a double point maps any plane not through the point to the point and any plane through the point to 00.
Taking this into account, it is still possible to define the binary relation that a point and a plane are polar to each other with respect to a degenerate point-wise quadric if the point is mapped to the plane by the polarity and a plane and a point are polar to each other with respect to a degenerate plane-wise quadric if the plane is mapped to the point by the polarity.
2.2.1 Notation
We write the point-wise, resp., plane-wise aspect of a quadric as $\mathbf{{Q}}$, resp., $\mathbf{{Q}}^{*}$, and the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ for the complete quadric. When either is regular, then so is the other and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and they satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=0$. In discussions limited to regular quadrics, we mostly refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role we employ the full notation introduced here.
2.2.2 Ring contact
The condition analogous to double contact is ring contact: the quadrics
$B=A+cpp^{T}$ (8)
meet in the conic where $p$ intersects $A$ and they have the same tangent planes at all these shared points. We call $p$ the ring plane and
refer to its polar point as the ring points. By the same reasoning as in the conic case, we have
$BP=AP+cpp^{T}P=AP$ (9)
when $P$ is incident to $p$ and thus
Lemma 2.2.
The polarities on two quadrics in ring contact act the same when restricted to a mapping between the points of the ring plane and the planes of the ring points. Conversely, if two polarities coincide on a plane, then the corresponding quadrics are in ring contact with respect to this plane.
Remark.
In the case where ring point is incident to the ring plane, i.e. the ring plane is a tangent plane and the ring point is its point of contact, the conic of ring contact is a (doubled) pair of lines. This case can be distinguished as X-contact though we include it under the umbrella term “ring contact”.
2.2.3 Some contact lemmas
We conclude this section with a few lemmas that will be useful for the proof the 8-quadric theorem.
Lemma 2.3.
If 3 quadrics are mutually in ring contact and at least one is regular, then all 3 are in the same ring contact pencil, i.e., there are no ring contact triangles.
Proof.
If $A,B$ and $C$ are in ring contact then there exist planes $p,q$ and $r$ such that
$B=A+pp^{T},\qquad C=B+qq^{T},\quad\mathrm{and}\quad\qquad A=\alpha C+rr^{T}$
where $\alpha$ is a scalar (note that we are not entitled to arbitrarily scale it to $1$ because we already fixed the relative scaling of $A$ and $C$).
Then
$C=A+pp^{T}+qq^{T}$
implying that
$(1-\alpha)C=pp^{T}+qq^{T}+rr^{T}.$
However, the right hand side has rank at most 3: it is either a cone with the intersection point of $p,q$ and $r$ as its vertex, or they meet in an axis through which it is a plane pair, or all three planes coincide. But we may, without loss of generality, choose $C$ to be the regular quadric and thus have rank $4$, implying that $\alpha=1$. Thus the left side vanishes and then, because 3 double planes may only be dependent if they coincide, $p,q$ and $r$ are necessarily the same plane. ∎
We remark that the same cannot be said of conics in double contact since rank 3 conics are nondegenerate, i.e., double contact triangles can exist, for instance
$A=\begin{pmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\qquad B=\begin{pmatrix}-1&0&0\\ 0&2&0\\ 0&0&2\end{pmatrix},\qquad C=\begin{pmatrix}-1&0&0\\ 0&2&0\\ 0&0&1\end{pmatrix}.$
On the other hand, if the chords of contact $p,q,r$ all meet in a point, then the span of $p^{2},q^{2},r^{2}$ contains only line pairs through this common point so we have the following lemma:
Lemma 2.4.
If 3 conics are mutually in double contact with concurrent chords of contact, then all 3 are in the same double contact pencil.
3 The main theorems
Theorem 3.1 (8-conic theorem).
In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). This conic is unique if not all the common points of condition (ii) are the same.
Theorem 3.2 (8-quadric theorem).
In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists an eighth quadric such that the completed cube satisfies (i) and (ii). This quadric is unique if not all the common axes of condition (ii) are the same.
3.1 Notation and terminology
The 8-conic and 8-quadric theorems are concerned with assigning conics (or quadrics) to (some or all of) the vertices of a combinatorial cube graph so that they satisfy certain edge and face relations. We introduce common terminology and notation for these configurations.
Definition 3.1.
A Penrose $n$-configuration is an assignment of $n$ conics (quadrics) to the vertices of a cube. The configuration is regular if the conics (quadrics) are regular.
Both the 8-conic and 8-quadric theorems assume that a regular Penrose 7-configuration is given and assert a unique completion to an 8-configuration. Note that this eighth conic (or quadric) is not necessarily unique or regular (see Thm. 3.3).
We can immediately count that a Penrose 7-configuration has 4 completed vertices, 9 completed edges, and 3 completed faces.
Next, we introduce notation for the vertices and faces of the cube. Consult Fig. 5, left. We decompose the cube vertices into two interlocked tetrahedra, and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $F^{ij}_{km}$ where $ijkm$ is a permutation of $0123$.
In condition (ii) of the 8-conic theorem, we call the common point of the chords of contact on a face the face point of the face; for condition (ii) of the 8-quadric theorem we call the common axis of the ring planes the face axis of the face.
Furthermore, in what follows we assume that the empty vertex of a Penrose 7-configuration is $\mathbf{{T}}^{0}$.
Figure 5: The cube graph of the 8-conic theorem (left) and an example of 4 conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2},\mathbf{{T}}^{1},\mathbf{{T}}^{3}$ satisfying condition (ii) (right).
The face figure of a Penrose conic configuration. Consider two conics corresponding to opposite nodes in a face of the cube, e.g. $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+pp^{T}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+qq^{T}$ with $p,q$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Hence $\mathbf{{T}}^{1}-\mathbf{{T}}^{3}=pp^{T}-qq^{T}=\frac{1}{2}(de^{T}+ed^{T})$ with $d=p+q$ and $e=p-q$ being a line pair of common chords in the pencil spanned by $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Thus the chords of contact meet at the point of intersection of a pair of common chords $e,d$. Since there are at most 3 such pairs of common chords the set of conics being in double contact with $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$ splits into at most 3 families, such that condition (ii) holds automatically for all the conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can explicitly be written in terms of $d,e$ and $\mathbf{{S}}_{0}$ (see Figure 5 and [Sal17, p.262]).
3.2 Penrose’s 8-conic theorem
The connection of the 8-conic theorem and the 8-quadric theorem is based on the following observation: Given two regular quadrics in ring contact and a plane that is not tangent to either quadric, slicing the quadrics with the plane yields two regular conics in double contact. In this slice, the common axis of the ring planes is cut in the common point of the chords of contact of the conics. Conversely, if two regular conics in double contact are given, it is possible to extrude them to obtain two regular quadrics in ring contact. (Similar remarks apply to the plane-wise view.)
The extrusion needed for the 7-configurations encountered in the following proofs is provided by the following lemma, whose proof appears below in Sec. 4.
Lemma 3.1 (Extrusion).
Given a regular Penrose 7-conic configuration, it is possible to extrude it to a regular Penrose 7-quadric configuration.
Proof of Penrose-7 conic theorem.
Apply Lemma 3.1 to the given regular Penrose 7-conic configuration to obtain an extrusion to a regular Penrose 7-quadric configuration. Then apply Thm. 3.2 to complete this to a Penrose 8-quadric configuration. The intersection of the plane of the original 7-conic configuration with the complete 8-quadric configuration is the desired regular Penrose 8-conic configuration. Face points of the conic configuration are distinct exactly when the corresponding face axes of the quadric configuration are distinct. Hence the eighth conic is unique exactly when the face points are not all the same. ∎
Remark. Due to factors explained in the next section, the proof presented below for the 8-quadric theorem does not cover the case that all the face axes, resp., points, are identical, hence the 8-conic theorem proof above does not cover the case that all chords of contact of the configuration pass through a single point.
3.3 Penrose’s 8-quadric theorem
We begin with three lemmas.
Let $i$ and $j\in\{0,1,2,3\}$, $m$ and $n$ in $\{1,2,3\}$ and $P^{m}_{i}$ denote the ring point for $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{m}$, $p^{m}_{i}$ denote the corresponding ring plane, $\Lambda^{mn}_{ij}$ the common axis of the ring planes for the face $F^{mn}_{ij}$ and $\lambda^{mn}_{ij}$ the common spear (point-wise line) of the ring points. We call these the face axes, resp., the face spears. By assumption, they are well defined. Equivalently, given any one of the three complete faces of the 7-configuration, its four ring planes are not all identical. Consider, for instance, the face $F^{12}_{03}$. By Lemma 2.3, we immediately conclude that $\mathbf{{S}}_{0}$ is not in ring contact with $\mathbf{{S}}_{3}$ and $\mathbf{{T}}^{1}$ is not in ring contact with $\mathbf{{T}}^{2}$. By similar reasoning applied to the other two complete faces, we get the following lemma:
Lemma 3.2.
$\mathbf{{S}}_{i}$ is not in ring contact to $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{m}$ is not in ring contact to $\mathbf{{T}}^{n}$ for any of the seven quadrics of a regular Penrose 7-configuration.
Since a face axis is determined by any two ring planes on a face, and every face of the 7-configuration has at least two ring planes, the axes and spears of the incomplete faces are also uniquely determined. In particular, the shared axis of the ring planes for the incomplete face $F_{ij}^{0n}$ must go through the ring planes for the two edges already given, so we set
$\Lambda^{0n}_{ij}:=p^{n}_{i}\wedge p^{n}_{j}$ (10)
and dually
$\lambda^{0n}_{ij}:=P^{n}_{i}\vee P^{n}_{j}.$ (11)
These will both be nonzero since otherwise, if we had $P^{n}_{i}=P^{n}_{j}$ or $p^{n}_{i}=p^{n}_{j}$, $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ would have to be in ring contact along the common conic of intersection with $\mathbf{{T}}^{n}$, contradicting Lemma 3.2.
We now demonstrate a global condition satisfied by a regular quadric 7-configuration which has no analogy in the case of conics.
Lemma 3.3 (Concurrent axes and spears).
For a regular Penrose 7-quadric configuration: If two of the face axes at $\mathbf{{S}}_{0}$ are distinct, then at each complete vertex, all three face axes are distinct. In this case, all ring planes and face axes are concurrent in a point $O$. Otherwise there is a single face axis $\Omega$ common to all ring planes. Dually, all ring points and face spears are concurrent in a plane $o$ or there is a single face spear $\omega$ common to all ring points. $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
Proof.
Assume $\Lambda^{23}_{01}=\Lambda^{13}_{02}$. Since each axis contains two of the ring planes at $\mathbf{{S}}_{0}$, all three of these ring planes pass through this axis so that it is also identical to $\Lambda^{12}_{03}$.
This contradicts the assumption that two of the three face axes $\Lambda^{ij}_{0k}$ are distinct. Hence all three are distinct if any two are.
We show now that the $\Lambda^{ij}_{0k}$ cannot all lie in a plane. Indeed, if they do, then that plane must be identical to the three ring planes $p^{j}_{0}$ but by the regularity of $\mathbf{{S}}_{0}$ this would imply that $\mathbf{{T}}^{i}$ is in ring contact to $\mathbf{{T}}^{j}$ for $i,j=1,2,3$ contradicting Lemma 3.2.
Hence the three axes at $\mathbf{{S}}_{0}$ are distinct and do not lie in a plane, and the ring planes are therefore distinct, and have a common point $O$, which is therefore also a common point of the axes.
Apply the same argument to the other complete vertices $\mathbf{{T}}^{i}$ to conclude that the axes of the faces meeting at $\mathbf{{T}}^{i}$ have a common point $O^{\prime}$. We show that $O^{\prime}=O$. Indeed, two of these axes belong to $\{\Lambda^{ij}_{0k}\}$, that were shown above to meet in $O$. So $O^{\prime}=O$. Since this applies to all 3 incomplete faces, all 6 axes are concurrent in $O$.
We now consider the case that two of the axes at $\mathbf{{S}}_{0}$ are the same. Then by the above, all three axes $\Lambda^{ij}_{0k}$ are the same, call it $\Omega$. By reasoning similar to the above, we can propagate this condition to the three remaining face axes, to show they also have to be equal to $\Omega$, and consequently that all ring planes pass through $\Omega$.
By regularity and duality the analogous result can be obtained for the ring points and the face spears. Furthermore, by Lemma 2.2, $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
We will also need this standard geometric result (see e.g. [Zie12, §4.3]):
Lemma 3.4 (Conical Octahedron Lemma).
Three conics in $\mathbb{R}P^{3}$, without a common point, that intersect pairwise in two points (including the possibility that the two points coincide and the conics thus share a tangent line), lie on a unique quadric.
Note that the conics cannot lie in a plane since two such conics have 4 points in common. The reader is also referred to the “coordinate version” of the 8-quadric proof below, which provides an independent proof of the lemma.
3.3.1 The proof
We are now in a position to prove the 8-quadric theorem, subject to the following restrictions.
First, we assume that two of the face axes at $\mathbf{{S}}_{0}$ are distinct. By Lemma 3.3, the alternative is that all six face axes are the same. This case will be handled using other methods in [ACG+].
We also assume for the moment that $o$ is not incident to $O$. This is equivalent to assuming the 7 quadrics of the configuration do not all share a common contact element. We postpone to Sec. 4.1 the proof that this condition is not necessary, since it depends on elements of the proof of Lemma 3.1.
Figure 6 depicts the completed cube configuration which we must construct. There are 6 axes corresponding to condition (ii), one for each face. Each face axis is the common line of the 4 ring planes on the edges of the face. As mentioned above, we restrict our attention to the first case in Lemma 3.3, whereby the 3 axes of the complete faces of the 7-configuration are not all identical and have a common point $O$, as shown in Figure 6. The dual plane $o$ is taken as the plane at infinity in this figure, thus the cones of ring contact are all cylinders through the conics $\mathbf{{R}}^{k}_{j}$ where $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{k}$ make ring contact. The 4 ring points on a face lie on six distinct spears in $o$.
Proof of the Penrose 8-quadric theorem.
Under the assumption that two of the faces axes at $\mathbf{{S}}_{0}$ are distinct, it suffices to construct a unique quadric $\mathbf{{T}}^{0}$ such that: for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $F^{0j}_{km}$ have a common axis.
Figure 6: Eight quadrics in ring contact and the structure of all 12 contact rings $\mathbf{{R}}_{i}^{k}$ together with the 6 axes $\Lambda_{ij}^{kn}$, sharing the point $O$. On each axis the four rings corresponding to a face of the cube diagram meet.
To begin with, we consider the planes and points that must necessarily be the ring planes and ring points for such a quadric to complete the cube. The ring plane for $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ will have to go through these axes for both faces adjacent to the desired edge representing ring contact:
$p^{0}_{i}:=\Lambda^{0n}_{ij}\vee\Lambda^{0j}_{in}$ (12)
(note: the joining plane of the two lines is defined because they meet in a point) and dually
$P^{0}_{i}:=\lambda^{0n}_{ij}\wedge\lambda^{0j}_{in}.$ (13)
To see that neither of these vanish, assume the contrary, that $\Lambda^{0n}_{ij}=\Lambda^{0j}_{in}$ (or its dual). Then by definition (10), $\Lambda^{0n}_{ij}$ would be a shared line common to all four of the planes $p_{i}^{n}$, $p_{j}^{n}$, $p_{n}^{j}$ and $p_{i}^{j}$ which would then at the same time coincide with $\Lambda^{ij}_{0n}$, $\Lambda^{jn}_{i0}$ and $\Lambda^{in}_{0j}$, contradicting the assumption that the latter are not all identical.
[…]
We first dispose of case ()(--)( - - ). Here, the shared plane of the $\Lambda^{0n}_{ij}$, doubled, is the final point-wise quadric and the shared common point of the $\lambda^{0n}_{ij}$, doubled, is the final quadric plane-wise. We now show that this point and plane are incident.
Claim: In case ()(--)( - - ), the common point $P=P_{1}^{0}=P_{2}^{0}=P_{3}^{0}$ of the $\Lambda^{ij}_{0n}$ is incident to the common plane $p=p_{1}^{0}=p_{2}^{0}=p_{3}^{0}$ of the $\lambda^{ij}_{0n}$, therefore $p^{2}$ (and plane-wise, $P^{2}$) is in $X$ contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2}$ and $\mathbf{{S}}_{3}$.
Proof of claim.
This would be shown if p were tangent to any of $\mathbf{{S}}_{1},\mathbf{{S}}_{2}$ and $\mathbf{{S}}_{3}$ because $P$ is polar to $p$ for all 3 of them and so if $p$ were tangent to $\mathbf{{S}}_{j}$, $P$ would be the point of tangency and thus incident to $p$. Therefore, for the sake of contradiction, we assume that $p$ is not incident to $P$ which implies that $\mathbf{{R}}_{j}=p\cap\mathbf{{S}}_{j}$ is a nondegenerate conic.
Then $\mathbf{{R}}_{i}$ and $\mathbf{{R}}_{j}$ are in double contact with $\Lambda^{0n}_{ij}$ as chord of contact because
$\Lambda^{0n}_{ij}\cap\mathbf{{R}}_{i}=\Lambda^{0n}_{ij}\cap\mathbf{{S}}_{i}=p^{n}_{j}\cap p^{n}_{i}\cap\mathbf{{S}}_{i}=p^{n}_{j}\cap p^{n}_{i}\cap\mathbf{{T}}^{n}=\Lambda^{0n}_{ij}\cap\mathbf{{T}}^{n}=p^{n}_{i}\cap p^{n}_{j}\cap\mathbf{{S}}_{j}=\Lambda^{0n}_{ij}\cap\mathbf{{R}}_{j}$ (14)
The first equality follows from the fact that the intersection points of $\mathbf{{S}}_{i}$ with $\Lambda^{0n}_{ij}$ necessarily occur on $\mathbf{{R}}_{i}$ since as point sets $\Lambda^{0n}_{ij}\subset p^{0}_{i}$, the second equality follows from the definition (10) of $\Lambda^{0n}_{ij}$, the third from the ring contact of $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{n}$ in the plane $p^{n}_{i}$ and the rest of the equalities follow by symmetrical reasoning. Both $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ are in contact with $\mathbf{{T}}^{n}$ at these intersection points, so slicing the common tangent planes with $p$ yields the lines of double contact of $\mathbf{{R}}_{i}$ and $\mathbf{{R}}_{j}$. Therefore, under the assumption that $P$ and $p$ are not incident, we obtain the conclusion that the three conics $\mathbf{{R}}_{i}$ are mutually in double contact with chords of contact $\Lambda^{0n}_{ij}$, however these are concurrent in the point $O$, so by Lemma 2.4 must be in the same double contact pencil which would imply that the 3 axes $\Lambda^{0n}_{ij}$ coincide which we’ve already seen cannot be the case. Therefore $P$ is incident to $p$. ∎
The remainder of the proof assumes that the 3 axes are linearly independent (that is, the cases (++)(++)( + + ) and (+)(+-)( + - ). Dualizing the arguments then takes care of the remaining case (+)(-+)( - + ).)
When the 3 axes are linearly independent, the rings of contact are
$\mathbf{{R}}^{0}_{i}:=\mathbf{{S}}_{i}\cap p^{0}_{i}.$ (15)
The conics $\mathbf{{R}}^{0}_{i}$ and $\mathbf{{R}}^{0}_{j}$ either meet in two points on $\Lambda^{0n}_{ij}$ or $\Lambda^{0n}_{ij}$ is a shared tangent line, because $\Lambda^{0n}_{ij}$ is the intersection line of their respective planes and
$\Lambda^{0n}_{ij}\cap\mathbf{{R}}^{0}_{i}=\Lambda^{0n}_{ij}\cap\mathbf{{R}}^{0}_{j}$
by the same reasoning as in equation (14).
To complete the cube diagram and prove the theorem, we apply Lemma 3.4 to obtain a unique quadric containing all three conics that is in ring contact with $\mathbf{{S}}_{i}$ along $\mathbf{{R}}^{0}_{i}$ with ring point $P^{0}_{i}$ for $i=1,2,3$.
Using this fact, we now provide a synthetic proof of the 8-quadric theorem, followed by a self-contained coordinate-based proof which includes a proof of Lemma (3.4) as part of the construction.
Recall that to complete the proof we must show that for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $F^{0j}_{km}$ have a common axis.
Condition (ii) is trivially fulfilled since we have used the face axes $\Lambda^{0j}_{km}$ to construct the new ring planes.
We now show condition (i) is also fulfilled. The polarity of $\mathbf{{R}}^{0}_{i}$ in the plane $p^{0}_{i}$ maps a point $P$ to the line of intersection of $p^{0}_{i}$ and the polar plane of $P$ with respect to any quadric containing $\mathbf{{R}}^{0}_{i}$: $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ in particular. By equations (10) and (11), and Lemma 2.2, the polarities on $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$ and $\mathbf{{T}}^{0}$ all behave the same way when restricted to points on $\Lambda^{0n}_{jk}$ and planes through $\lambda^{0n}_{jk}$. Combining these two facts and applying them to the cases $i=j$ and $i=k$, any point on $\Lambda^{0n}_{jk}$ is mapped under the respective polarities of $\mathbf{{R}}^{0}_{j}$ and $\mathbf{{R}}^{0}_{k}$ to two lines which together span the polar plane of $P$ with respect to any of the quadrics $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$, $\mathbf{{T}}^{0}$. Finally, by considering the same scenario with $\Lambda^{0k}_{jn}$ in place of $\Lambda^{0n}_{jk}$, and appealing to the definitions of $P^{0}_{j}$ and $p^{0}_{j}$ in equations (12) and (13), we conclude that the polarities on $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{0}$ agree when restricted to planes through $P^{0}_{j}$ and points on $p^{0}_{j}$ so $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{j}$ are in ring contact by Lemma 2.2. ∎
Coordinate-based proof of 8-quadric theorem.
Here we construct the point-wise matrix of the eighth quadric in terms of the point-wise matrices of the seven given ones. Dualizing yields the plane-wise matrix.
Let the matrices representing the quadrics $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{m}$ in the ordered basis $(O,X_{1},X_{2},X_{3})$ where
$X_{n}=\Lambda^{0n}_{ij}\wedge o.$ (16)
That this forms a basis follows from the independence of the $\Lambda^{0n}_{ij}$ since we are working in case (I), together with the fact that $O$ is not incident to $o$. Because $o$ is polar to $O$ for all of the given quadrics, and $X_{1},X_{2}$ and $X_{3}$ span $o$, their representing matrices in this basis may be chosen to have the form
$\begin{pmatrix}1&0&0&0\\ 0&\ast&\ast&\ast\\ 0&\ast&\ast&\ast\\ 0&\ast&\ast&\ast\end{pmatrix}.$ (17)
Let $\mathbf{{S}}_{1}$ be represented by
$A=\begin{pmatrix}1&0&0&0\\ 0&a_{11}&a_{12}&a_{13}\\ 0&a_{12}&a_{22}&a_{23}\\ 0&a_{13}&a_{23}&a_{33}\end{pmatrix},$ (18)
$\mathbf{{S}}_{2}$ by
$B=\begin{pmatrix}1&0&0&0\\ 0&b_{11}&b_{12}&b_{13}\\ 0&b_{12}&b_{22}&b_{23}\\ 0&b_{13}&b_{23}&b_{33}\end{pmatrix},$ (19)
and $\mathbf{{S}}_{3}$ by
$C=\begin{pmatrix}1&0&0&0\\ 0&c_{11}&c_{12}&c_{13}\\ 0&c_{12}&c_{22}&c_{23}\\ 0&c_{13}&c_{23}&c_{33}\end{pmatrix}.$ (20)
Then $\mathbf{{R}}^{0}_{1}$ is in the plane $p^{0}_{1}$ spanned by $O,X_{2}$ and $X_{3}$ and is the restriction of $\mathbf{{S}}_{1}$ thereto and thus can be represented in the ordered basis $(O,X_{2},X_{3})$ by the matrix
$\bar{A}=\begin{pmatrix}1&0&0\\ 0&a_{22}&a_{23}\\ 0&a_{23}&a_{33}\end{pmatrix}.$ (21)
The conics $\mathbf{{R}}^{0}_{2}$ and $\mathbf{{R}}^{0}_{3}$ can similarly be represented in the bases $(O,X_{1},X_{3})$ and $(O,X_{1},X_{2})$ by
$\bar{B}=\begin{pmatrix}1&0&0\\ 0&b_{11}&b_{13}\\ 0&b_{13}&b_{33}\end{pmatrix}\qquad\mathrm{and}\qquad\bar{C}=\begin{pmatrix}1&0&0\\ 0&c_{11}&c_{12}\\ 0&c_{12}&c_{22}\end{pmatrix}$ (22)
respectively. If $\mathbf{{R}}^{0}_{2}$ and $\mathbf{{R}}^{0}_{3}$ are tangent then $b_{11}=c_{11}=0$, otherwise
$\begin{pmatrix}1&0\\ 0&c_{11}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&b_{11}\end{pmatrix}$ (23)
represents the pair of intersection points in the basis $(O,X_{1})$ for $\Lambda^{01}_{23}$. This implies in particular that
$c_{11}=b_{11}$ (24)
and by similar reasoning
$a_{22}=c_{22}\qquad\mathrm{and}\qquad a_{33}=b_{33}.$ (25)
Therefore
$D=\begin{pmatrix}1&0&0&0\\ 0&d_{11}&c_{12}&b_{13}\\ 0&c_{12}&d_{22}&a_{23}\\ 0&b_{13}&a_{23}&d_{33}\end{pmatrix},$ (26)
where
$d_{11}:=c_{11}=b_{11},\qquad d_{22}:=a_{22}=c_{22}\qquad\mathrm{and}\qquad d_{33}:=a_{33}=b_{33},$ (27)
represents a quadric $T_{0}$ containing the three conics $\mathbf{{R}}^{0}_{1},\mathbf{{R}}^{0}_{2}$ and $\mathbf{{R}}^{0}_{3}$. To show that $\mathbf{{T}}^{0}$ satisfies the required ring contacts as well as conditions (ii) and (ii)’, it is sufficient to show that it has $\Lambda^{0n}_{ij}$ polar to $\lambda^{0n}_{ij}$ for all $\{i,j,n\}=\{1,2,3\}$ because then the intersection
$P^{0}_{i}:=\lambda^{0n}_{ij}\wedge\lambda^{0j}_{in}$ (28)
is polar to $p^{0}_{i}$ with respect to both $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ and is thus the ring point corresponding to the ring of contact $\mathbf{{R}}^{0}_{i}$.
Because $\mathbf{{S}}_{1}$ and $\mathbf{{S}}_{2}$, by virtue of ring contact with $\mathbf{{T}}^{3}$, both have $\Lambda^{03}_{12}$ polar to $\lambda^{03}_{12}$, both have $X_{3}$ polar to $x_{3}$ where
$x_{n}=\lambda^{0n}_{ij}\vee O.$ (29)
Therefore, because the fourth column of each matrix gives the plane coordinates of the image of $X_{3}$:
$a_{13}:a_{23}:a_{33}=b_{13}:b_{23}:b_{33}.$ (30)
By similar reasoning,
$a_{12}:a_{22}:a_{23}=c_{12}:c_{22}:c_{23}$ (31)
and
$b_{11}:b_{12}:b_{13}=c_{11}:c_{12}:c_{13}.$ (32)
This implies that
$b_{13}:a_{23}:d_{33}=b_{13}:b_{23}:b_{33},$ (33)
$a_{12}:a_{22}:a_{23}=c_{12}:d_{22}:a_{23}$ (34)
and
$d_{11}:c_{12}:b_{13}=c_{11}:c_{12}:c_{13}.$ (35)
Therefore, $\mathbf{{T}}^{0}$ has $X_{n}$ polar to $x_{n}$ for $n=1,2,3$ completing the proof. ∎
3.3.2 When the eighth quadric is not regular
The case (++)(++)( + + ), in which both triples of axes and spears are linearly independent, clearly produces regular matrices, hence the eighth quadric is regular. This section considers what kinds of degenerate quadrics arise in the other cases. We remark that by equation (29) the $\lambda^{0n}_{ij}$ meet in a point exactly when the $x_{n}$ share a common line, i.e., when the resulting matrix $D$ is degenerate. It must, furthermore, have rank $3$ because the $\lambda^{0n}_{ij}$ do not coincide and $o$ is not incident to $O$ and thus is independent of the $x_{n}$. Therefore in case (+)(+-)( + - ), if the $\lambda^{0n}_{ij}$ are dependent, $\mathbf{{T}}^{0}$ is a cone. In case (+)(-+)( - + ), we dualize and obtain a conic. The final point-wise quadric in ring contact is the plane of this conic, doubled. In case ()(--)( - - ), the final point-wise quadric is again a double plane but now it does not arise as the plane of a conic but rather as a contact element together with a pair of lines: the lines of $X$-contact (note however that the $\mathbf{{S}}_{j}$ are not in $X$-contact with each other: as discussed in, e.g. [BO20], there are several different contact relations for which two quadrics may meet in a pair of lines, $X$-contact being the special case where this is the entirety of the intersection; in the present case, this will hold for each $\mathbf{{S}}_{j}$ with the double plane $\mathbf{{T}}^{0}$ but not for the $\mathbf{{S}}_{j}$ with each other). From this it is clear that the full information of the quadrics is not contained in the point-wise or plane-wise description alone but requires knowledge of both (as well as possibly lines). However, in the present paper we leave the precise formulation somewhat one-sided, leaving a more principled exposition to future work.
3.3.3 Refinement of Penrose 8-quadric theorem
The above discussion contains a proof of the following refinement of Thm. 3.2.
Theorem 3.3.
Given a regular Penrose 7-quadric configuration. Let $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ be the quadric that, by Thm. 3.2, completes the configuration. Consider the 3 face axes and 3 face spears of the cube faces meeting at the eighth vertex. Then:
  1. 1.
    If the axes do not lie in a plane and the spears do not pass through a point, $\mathbf{{Q}}$ and $\mathbf{{Q}}^{*}$ are regular.
  2. 2.
    If the axes lie in a plane and the spears do not pass through a point, $\mathbf{{Q}}$ is a double plane and $\mathbf{{Q}}^{*}$ is a conic,
  3. 3.
    If the axes do not lie in a plane and the spears pass through a point, $\mathbf{{Q}}$ is a cone and $\mathbf{{Q}}^{*}$ is a double point,
  4. 4.
    If the axes lie in a plane and the spears pass through a point, $\mathbf{{Q}}$ is a double plane and $\mathbf{{Q}}^{*}$ is an incident double point.
There is an analogous version of this theorem for a Penrose 7-conic configuration. Out of space considerations, we omit the details.
4 Proof of Extrusion Lemma
We now turn to the promised proof of Lemma 3.1: Given a regular Penrose 7-conic configuration, it is possible to extrude it to a regular Penrose 7-quadric configuration.
Proof of extrusion lemma.
Let $(P_{1},P_{2},P_{3})$ be an ordered basis of the plane containing a given conic configuration and $O$ a point not on this plane. Then if we choose a plane $o$ with homogeneous coordinates $[u_{0}:u_{1}:u_{2}:u_{3}]$ in the basis dual to $(O,P_{1},P_{2},P_{3})$, any conic in the plane, represented by a matrix
$A=\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{12}&a_{22}&a_{23}\\ a_{13}&a_{23}&a_{33}\end{pmatrix}$ (36)
in the basis $(P_{1},P_{2},P_{3})$ can be extruded to a quadric
$\bar{A}=\begin{pmatrix}u_{0}&u_{1}&u_{2}&u_{3}\\ u_{1}&a_{11}&a_{12}&a_{13}\\ u_{2}&a_{12}&a_{22}&a_{23}\\ u_{3}&a_{13}&a_{23}&a_{33}\end{pmatrix}$ (37)
going through the conic and having $O$ polar to $o$.
The quadric configuration resulting from this construction applied to the 7 conics of the hypothesis of the 8-conic theorem will automatically satisfy conditions (i) and (ii) of the 8-quadric theorem, provided that the chords of contact of condition (i) for the conic case are extruded to ring planes through $O$. That is, a regular Penrose 7-conic configuration produces a regular Penrose 7-quadric configuration.
Let $A$ and $B$ represent two conics in double contact. Then
$\mathrm{rank}(A+rB)=1$ (38)
for some scalar $r$ and $\bar{A}$ will be in ring contact with
$\bar{B}(t)=\begin{pmatrix}u_{0}&u_{1}&u_{2}&u_{3}\\ u_{1}&tb_{11}&tb_{12}&tb_{13}\\ u_{2}&tb_{12}&tb_{22}&tb_{23}\\ u_{3}&tb_{13}&tb_{23}&tb_{33}\end{pmatrix}$ (39)
with ring plane going through $O$ exactly when $t=r$. Thus in order to perform the extrusion, it suffices to simultaneously scale all the matrices representing the 7 conics so that for any pair in double contact represented by the matrices $A$ and $B$ we have
$\mathrm{rank}(A-B)=1.$ (40)
To begin with, the scaling of the matrix representing $\mathbf{{S}}_{0}$ can be chosen arbitrarily and then the matrices representing $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$, and $\mathbf{{T}}^{3}$ are uniquely defined by equation (40). Then it only remains to scale $\mathbf{{S}}_{1},\mathbf{{S}}_{2}$ and $\mathbf{{S}}_{3}$ so that equation (40) holds for both double contacts of each. Let $A,B$ and $C$ be the appropriately scaled matrices representing $\mathbf{{S}}_{0},\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{2}$ respectively:
$A-B=V^{1}_{0},\qquad A-C=V^{2}_{0}$ (41)
where $V^{j}_{i}$ is a rank 1 matrix representing the doubled chord of contact. Then let $D$ be the matrix representing $\mathbf{{S}}_{3}$ scaled so that
$D-B=V^{1}_{3}.$ (42)
Then all four matrices are scaled so the double contact of $\mathbf{{T}}^{2}$ and $\mathbf{{S}}_{3}$ in these scalings must be written
$D-rC=V^{2}_{3}$ (43)
for some determined scalar $r$. However, by sequentially eliminating $A,B$ and $D$ from equations (41)-(43), we obtain
$(r-1)C=V^{1}_{3}+V^{2}_{0}-V^{1}_{0}-V^{2}_{3}.$ (44)
Because of condition (ii) of the hypothesis of the 8-conic theorem, all the chords of contact for the face $(\mathbf{{S}}_{0},\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{S}}_{3})$ share a point whose representative in the basis $(P_{1},P_{2},P_{3})$ is necessarily in the common kernel of $V^{1}_{3}$, $V^{2}_{0}$ ,$V^{1}_{0}$ and $V^{2}_{3}$. However, by the regularity assumption on $\mathbf{{T}}^{2}$, $C$ has trivial kernel which is only consistent with equation (40) in the case $r=1$.
Similar reasoning applied to the other two faces yields the scaling needed to establish the desired result.
4.1 Proof of the 8-quadric theorem with a common contact element
We can now give the promised proof of the 8-quadric theorem when there is a common contact element of all the quadrics.
Proof with common contact element.
Given a configuration of conics in double contact, equation (39) exhausts all possible extensions of $B$ to a quadric polarizing $O$ and $o$ and having all of these double contacts as slices of ring contacts, with $O$ as the point of concurrency of the contact axes and thus $o$ as the common plane of the ring points. Therefore, to prove the case of the 8-quadric theorem where $o$ is incident to $O$, we apply the 8-conic theorem to a planar slice of the 7-quadric configuration and then extrude the resulting complete 8-conic via equation (39), choosing the parameter $t$ that extrudes the seven conics obtained from the slicing process back to the seven quadrics we started with. The eighth extruded quadric is then the desired conclusion of the 8-quadric theorem. ∎
5 Examples of 8-quadric configurations
The aim of this section is to illustrate the structure of the 8-quadric theorem by constructing several examples step by step and, by slicing them, to obtain all the famous 2D-porisms mentioned in the introduction. Though the 8-quadric theorem was stated for regular quadrics, much of the reasoning in section 3.3 can be traced through degenerate cases like the ones in this present section. We provide some indications of necessary modifications in the logic of the proof in order to apply to these examples, however, we leave the delineation of the general conditions for allowable degeneracies, and the complete enumeration of the possible cube configurations that may result, to a future work dedicated solely to that subject.
Example 5.1.
Consult Fig. 7. We start with a sphere (green) as quadric $\mathbf{{S}}_{0}$. Then we choose 3 planes $p_{0}^{i}$ intersecting the sphere in 3 (blue) rings $\mathbf{{R}}_{0}^{i}$ such that each pair of rings meet in a pair of points (white). They form a conical octahedron, featured above in Lemma 3.4. The corresponding cones touching the sphere along those rings are taken as quadrics $\mathbf{{T}}^{i}$ (yellow). Now for each pair of cones the intersection $\mathbf{{T}}^{i}\cap\mathbf{{T}}^{j}$ consists of two conics (red). These six conics meet three at a time in the 8 yellow points, the common intersection $\mathbf{{T}}^{1}\cap\mathbf{{T}}^{2}\cap\mathbf{{T}}^{3}$ of all the three cones. We call this configuration a conical cube, whereby the yellow points represent the vertices and each red conic represents a pair of opposite edges of the cube.
We will choose one conic from each of the three pairs as a degenerate quadric $\mathbf{{S}}_{k},\{i,j,k\}=\{1,2,3\}$. This situation can be pictured as a limit of thin ellipsoids in ring contact with $\mathbf{{T}}^{i},\mathbf{{T}}^{j}$ whose planes of ring contact $p_{k}^{i},p_{k}^{j}$ finally converge to a single plane, which contains the chosen conic of the intersection.
Figure 7: Three cones in ring contact with a sphere (left). The prototype of this situation are 3 cylinders with pairwise orthogonal axes (right). The 3 rings (blue) form a conical octahedron, whereas the 6 conics of the intersections (red) form a conical cube.
There are now two possibilities for such a triple of conics $\{\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}\}$, one leading to a regular $T_{0}$ and one not: Either the 3 conics meet in a point pair, or they form a conical octahedron. In Figure 7, left, the first possibility is shown by the three red conics that meet in the yellow point in the middle surrounded by the three blue ring conics (with a second meeting point on the opposite side of the sphere), while the second possibility is shown by the remaining three red conics, that surround the point in the middle and extend farther out along the 3 yellow cones and meet pairwise in two points.
In the latter case the final quadric $\mathbf{{T}}^{0}$ is regular (directly given by Lemma 3.4). Since each quadric $\mathbf{{S}}_{i}$ degenerates to a conic, all 3 rings $\mathbf{{R}}_{i}^{k}$ on it also collapse to this conic and each axis $\Lambda_{ij}^{0k}$ is the intersection of two ring planes. In the former case, on the other hand, the final quadric $\mathbf{{T}}^{0}$ is a point pair. This case, however, cannot as easily be subsumed into the logic of section 3.3 since the final 3 axes are identical: $\Lambda^{01}_{23}=\Lambda^{02}_{13}=\Lambda^{03}_{12}$. To picture this in terms of a limit, we consider a sequence of shrinking ellipsoids, the $\mathbf{{S}}_{i}$ getting flatter and $\mathbf{{T}}^{0}$ getting thinner and finally collapsing to a line segment connecting the two points, all planes through either point ultimately being tangent planes enveloping the ellipsoid in the limit.
Although these configurations are quite elementary in 3D their slices with a plane yield nontrivial theorems in 2D. In particular, consider the plane $p$ through the ring points of the 3 cones: It intersects the cones in 6 lines tangential to a circle being the intersection of the sphere $\mathbf{{S}}_{0}$ with this plane. In the case of three concurrent conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, they intersect the plane $p$ in 6 points which form a hexagon circumscribing the circle of intersection with the sphere. This leads to a Brianchon configuration. In the case of three non-concurrent conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ they intersect the plane $p$ in 6 points which split up into two triangles circumscribing the circle of intersection. This leads to a Poncelet configuration. A slice with a general plane gives the double contact theorem and the 4-conic theorem from [EMCT74]. Note that a single 3D-configuration contains Brianchon figures in 4 different ways such that the Brianchon diagonals form a complete 4-point.
Figure 8: Left: Three concurrent conics of the conical cube induce a hexagon circumscribing a circle in the plane $p$ through the ring points of the cones and thus proves the Brianchon theorem. Right: In the case of three non-concurrent conics the 6 points split into two triangles. This figure proves the Poncelet theorem.
Example 5.2.
In the next example we consider three spheres $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ such that no sphere is completely contained in another one. All spheres intersect the plane at infinity in the same imaginary circle $\mathbf{{S}}_{0}$, the absolute of Euclidean geometry. Thus each sphere $\mathbf{{T}}^{i}$ is in ring contact with $\mathbf{{S}}_{0}$, where the quadric $\mathbf{{S}}_{0}$ as an imaginary degenerate conic represents at the same time the rings $\mathbf{{R}}_{0}^{i}$. We could proceed as in the previous example by choosing as quadrics $\mathbf{{S}}_{k}$ the intersection of the spheres $\mathbf{{T}}^{i}\cap\mathbf{{S}}_{0}$, which are circles (besides the absolute imaginary circle at infinity) and by slicing with a plane through the centers of the 3 spheres we would obtain again a Brianchon-like theorem, which turns out to be the theorem of the radical center of 3 circles.
Instead we choose cones enveloping each pair of spheres. If the spheres are disjoint there are 6 such cones in total (dual to the 6 conics in Example 5.1) and again there are two cases for the choice of a triplet of cones. In one case the cones share 2 tangent planes which are also common tangent planes of the 3 spheres. See Figure 9, left. Thus the final quadric $\mathbf{{T}}^{0}$ is this pair of planes and the rings $\mathbf{{R}}_{i}^{0}$ are 3 line pairs. This 3D-configuration was already used by G. Monge (according to [DW92]p.154) in order to prove a theorem nowadays bearing his name: The three pairs of external tangents of three circles meet in three points on a line.
In the other case the cones do not admit common tangent planes. See Figure 9, right. Now the rings $\mathbf{{R}}_{i}^{0}$ are obtained by intersecting each cone $\mathbf{{S}}_{i}$ with the corresponding plane $p_{i}^{0}$ which is the join of a pair of vertical axes. The final quadric is again guarantied by lemma 3.4. A planar slice through the centers of the spheres shows that 3 pairs of internal tangents of 3 circles are circumscribed to a conic. This theorem seems to be known only for the case, where the final conic degenerates into a point pair, see [Bak36],[Sal17].
Analogous to the Example 5.1 there are 4 ways to assign a pair of common tangent planes to 3 spheres. The 6 ring points of the cones form a complete 4-line.
Figure 9: Left: Original proof of G. Monge of his theorem about the external tangents of three circles. Right: A theorem about the internal tangents
Example 5.3.
Consider 3 planes $p_{0}^{i}$ tangent to a quadric $\mathbf{{S}}_{0}$ at points $P_{0}^{i}$. In order to see the rings $\mathbf{{R}}_{0}^{i}$ as real objects we choose here a regulus as quadric $\mathbf{{S}}_{0}$, thus each ring is a line pair contained in the regulus. Consult Fig. 10. This is the case referred to as X-contact in section 2.2, whereby the ring point is incident to the ring plane. We want to picture these line pairs as ‘quadrics’ $\mathbf{{T}}^{i}$, although the situation is more complex. They are in fact double planes when considered point-wise and double points when considered plane-wise, but in an X-contact pencil we can identify them with the line pairs themselves. We can imagine these ‘quadrics’ as the limit of a sequence of reguli all containing the line pair $\mathbf{{S}}_{0}\cap p_{0}^{i}$, as they get flatter and flatter.
Each pair of line pairs $\mathbf{{T}}^{i},\mathbf{{T}}^{j}$ meet in a pair of points on $\mathbf{{S}}_{0}$ and are contained in a pair of tangent planes since $\mathbf{{T}}^{i},\mathbf{{T}}^{j}\subset\mathbf{{S}}_{0}$. We choose the plane pairs to be the quadrics $\mathbf{{S}}_{k}$. Then the eighth quadric is the double plane spanned by the three points $P_{0}^{1},P_{0}^{2},P_{0}^{3}$ we started with, and the rings $\mathbf{{R}}_{i}^{0}$ are the 3 sides of this triangle, each side considered as a double line, see Figure 10.
The configuration described above is well known. The 3 plane pairs $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ form a hexahedron which is inscribed around a regulus such that 6 of the edges (a skew hexagon) lie in the regulus. Thus a planar slice of this configuration yields a hexagon inscribed in a conic such that opposite sides of the hexagon (corresponding to the plane pairs $\mathbf{{S}}_{i}$) meet in 3 points which lie on a line (the intersection of the slicing plane with the plane $\mathbf{{T}}^{0}$), the Pascal line of the hexagon. This 3D proof of Pascal’s Theorem was already found in 1826 by G. Dandelin [Dan26].
Figure 10: The proof of Pascal’s Theorem using a hexahedron inscribed around a regulus. The Pascal axis is the intersection of the planar slice (bottom) with the ideal plane of the hexahedron (top).
Finally we will look at the same configuration but construct it in reverse order, starting with a double plane $\mathbf{{S}}_{0}$ followed by three plane pairs $\mathbf{{T}}^{i}$, each plane pair meeting in a line contained in $\mathbf{{S}}_{0}$. Each pair of plane pairs intersects in two line pairs. In this special case the conical cube of the 6 line pairs are the opposite edges of a usual projective cube. As in the examples before a triplet of chosen line pairs might be concurrent or not. In the first case the 3 line pairs obviously split into two triples of lines, belonging to opposite vertices in the cube, see Figure 11 left. A planar slice yields Desargues’ theorem. For the other choice, the 3 line pairs form a closed zigzag and the planar slice yields Pascal’s theorem. In both cases the axis in the theorem is the intersection with the plane $\mathbf{{S}}_{0}$.
Figure 11: If three pairs of opposite edges on a hexahedron (a projective cube) are chosen, such that one line from each pair meet in two opposite vertices (left), a planar slice splits into two perspective triangles, proving Desargues’ Theorem (middle). Otherwise the three edge pairs form a zigzag on the cube and yield a hexagon in a planar slice, proving Pascal’s Theorem (right).
The above examples show the following common structure: The three quadrics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ give rise to 6 degenerate quadrics either by pairwise intersections (Examples 5.1 and 5.3) or by pairwise envelopes (Example 5.2). Depending on the choice of such a triplet $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, one from each pair, the final quadric $\mathbf{{T}}^{0}$ is either regular or a pair of points (or planes). This is summarized in Figure 12.
Ex.1 Refer to caption Ex.2 Refer to caption Ex.3 Refer to caption
Figure 12: Graphs of the 8-quadric configurations of the above examples. The corresponding 2D-theorems obtained by appropriate slices are: Ex.1 Poncelet/Brianchon, Ex.2 Original/Monge, Ex.3 Pascal/Desargues
6 Related research and outlook for future work
We recently became aware of an independent discovery of the 8-conic theorem in [Eid13], Sec. 9.2.
The research leading to this article has also produced two companion articles which provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9D space of quadrics. The proofs have a similar structure to each other, but advantages of the 3D parameter-space approach over the 2D approach become clear in the final stages of the proofs.
This article and [ACG+] are the result of a collaboration with Roger Penrose that arose after the authors of [CGN] contacted him to report on their results. A collaboration ensued that led, among other results, to the successful reconstruction of the original 3D quadric proof of Roger Penrose in the two forms mentioned above.
Roger Penrose also told us that as an undergraduate he first recognized the cube structure of the theorem from examples in which most of the conic sections are degenerate. He then developed a method for gluing together several such cubes to generate new cubes that contain more regular conics. In this way, he finally arrived at the formulation of the general theorem. Furthermore, the possibility of gluing several cubes together is also interesting for obtaining further theorems. For example, this method can be used to extend Poncelet’s theorem for triangles to general polygons. Another example is Neville’s theorem about three ellipses that have a common focal point in pairs: Here, a first application of the eight-conic theorem can be used to show that all three ellipses must have double contact with another conic. This then serves as a starting point for further cubes, attached to the first one. We are currently working on examples of this type of iterated application of the theorem.
Regarding cube configurations with degenerate conics or quadrics, further conditions are needed to ensure the validity of the main theorems, presented in this article. This can be investigated in parameter space by looking at limits of regular configurations. It seems very worthwhile to us to further develop these techniques making use of the projective geometric structures of parameter space, as done in [CGN] and [ACG+].